OpenStudy (anonymous):

The perimeter of a rectangle is 90 ft. Find the dimensions of the rectangle if the ratio of the length to the width is 8: 7

OpenStudy (anonymous):

answers could be 30 ft by 15 ft 26 ft by 19 ft 24 ft by 21 ft 20 ft by 25 ft

OpenStudy (anonymous):

@Easyaspi314 help please.

OpenStudy (anonymous):

Like we did before, if the ratio is 8:7, then that can be expressed as 8x:7x. So the perimeter is 2(8x) + 2(7x) = 16x + 14x = 30x so 30x = 90 x = 3 So the dimensions is 8x or 8(3) = 24 7x = 7(4) = 28 So it is a 28 by 24. Understand?

OpenStudy (anonymous):

i think i do but thats not any of the answers

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

Sorry...I had 7(4)...supposed to be 7(3) = 21 becuase we had x = 3 So it is a 28 by 21

OpenStudy (anonymous):

24 by 21

OpenStudy (anonymous):

again...8x by 7x is 24 by 21 (when x = 3)

OpenStudy (anonymous):

im sorry iv just had some trouble with this thanks for your help @Easyaspi314